# Black hole event horizons are cosmologically coupled

### v1

@@ Abstract
It is shown that an exactly static black hole event horizon cannot be embedded in a time-dependent geometry. Forcing it to do so results in a naked null singularity at the would-be horizon. Therefore, since the universe is expanding, black holes must couple to the cosmological expansion, which was suggested as the growth mechanism for supermassive black holes in galaxies, with implications for the dark energy puzzle.
@@ I Introduction
A major problem of contemporary astrophysics consists of the fact that supermassive black holes at the centres of galaxies must have grown very fast, which is problematic if one invokes standard astrophysical channels [1, 2]. Moreover, LIGO detections of gravitational waves from binary mergers involving black holes highlight the presence of stellar mass black holes in the mass gap where they should not exist according to well-established models of stellar evolution [3, 4, 5]. These astrophysical puzzles can potentially be resolved if black holes interact with the universe in which they are embedded and their masses evolve over cosmological time scales (“cosmological coupling”) [6].
Realistic black holes are not asymptotically flat but are embedded in the universe. The study of astrophysical black holes usually neglects the cosmic expansion due to the vast separation between astrophysical (local) scales and the cosmological scale. However, over billions of years, tiny interactions between black hole and surrounding universe, if present, accumulate and can become important. The recent Ref. [6] reports the observation of cosmological coupling between the masses of supermassive black holes and the scale factor of the Friedmann-Lemaître-Robertson-Walker (FLRW) universe, obtained from data of black hole populations in red elliptical galaxies in the redshift range $0<z\leq 2.5$. Furthermore, if black holes are non-singular objects with de Sitter-like macroscopic cores, they could even explain dark energy, which would effectively be relegated to the interiors of black holes [6], a phenomenology possibly extending to non-singular black hole mimickers, which are now the subject of a vast theoretical literature (e.g., [7, 8, 9]). Cosmological coupling is also in agreement with the recent DESI results on the redshift evolution of the dark energy equation of state [10]. The 2023 tentative observation of cosmological coupling [6] caused a lively debate in both the observational and theoretical communities during the past year, as the arguments supporting it seemed unconvincing to some of its critics-here we provide more convincing ones.
@@ II Static black hole event horizon in a time-dependent metric
as can be read off the angular part of the line element (2.1). The equation locating the apparent horizons [^2] This equation generalizes the well known relation locating the Schwarzschild event horizon $g^{RR}=\nabla^{c}R\nabla_{c}R=1-2m/R=0$ to any spherical geometry. (if they exist) is $\nabla^{c}R\nabla_{c}R=0$. A single root corresponds to a black hole or white hole horizon, a double root to a wormhole [16]. This equation locates also event horizons, and gives
@@
The two-dimensional metric induced on the surface $R=$ const. is static but the time-dependent geometry just outside the horizon has $T=T(t,r)$, $a=a(t,r)$ and the areal radius $R(t,r)=a(t,r)r$ varies in time and cannot match $R=$ const. on the would-be horizon. Therefore, the junction conditions [23, 24] cannot be satisfied there: the discontinuity in the metric $g_{ab}$ induced on this 2-surface cannot be handled with distributional techniques and the jump in the metric $g_{ab}$, unlike a jump in its transversal derivative (interpreted as due to a matter layer on a null shell) has no physical interpretation [24].
@@ III Radial geodesics and Ricci scalar at the horizon
Let us examine outgoing and ingoing radial null geodesics. We will show that they cannot be continued inside the would-be static black hole horizon, which is unphysical. Furthermore, the Ricci scalar of the geometry (2.1) diverges there. The difficulties arise from assuming an exactly static horizon in a time-dependent geometry, showing that this assumption is unphysical.
@@ III.1 Radial null geodesics
III.2 Ricci scalar
@@ Acknowledgements.
We are grateful to Kevin Croker for discussions. V.F. is supported by the Natural Sciences & Engineering Research Council of Canada (grant 2023-03234).

### v2

@@ Abstract
It is shown that an exactly static and spherically symmetric black hole event horizon cannot be embedded in a time-dependent geometry. Forcing it to do so results in a naked null singularity at the would-be horizon. Therefore, since the universe is expanding, black holes must couple to the cosmological expansion, which was suggested as the growth mechanism for supermassive black holes in galaxies, with implications for the dark energy puzzle.
@@ I Introduction
A major problem of contemporary astrophysics consists of the fact that supermassive black holes at the centres of galaxies must have grown very fast at redshifts $z>6$, which is problematic if one invokes standard astrophysical channels [1, 2]. Moreover, LIGO detections of gravitational waves from binary mergers involving black holes highlight the presence of stellar mass black holes in the mass gap where they should not exist according to well-established models of stellar evolution [3, 4, 5], which is another potential problem if confirmed. These astrophysical puzzles can potentially be resolved if black holes interact with the universe in which they are embedded and their masses evolve over cosmological time scales (“cosmological coupling”) [6] 111 To be specific, in this paper “cosmological coupling” refers to any evolution of a black hole event horizon with time or with the scale factor of the FLRW in which they are embedded. .
These astrophysical black holes are usually modelled as asymptotically flat Schwarzschild or Kerr black holes. However, realistic black holes are not asymptotically flat but are embedded in the universe. The study of astrophysical black holes usually neglects the cosmic expansion due to the vast separation between astrophysical (local) scales and the cosmological scale. However, over billions of years, tiny interactions between black hole and surrounding universe, if present, accumulate and can become important. The recent Ref. [6] reports plausible evidence for cosmological coupling between the masses of supermassive black holes and the scale factor of the Friedmann-Lemaître-Robertson-Walker (FLRW) universe, obtained from data of black hole populations in red elliptical galaxies in the redshift range $0<z\leq 2.5$. Furthermore, if black holes are non-singular objects with de Sitter-like macroscopic cores, they could even explain dark energy, which would effectively be relegated to the interiors of black holes [6], a phenomenology possibly extending to non-singular black hole mimickers, which are now the subject of a vast theoretical literature (e.g., [7, 8, 9, 10, 11, 13, 14, 15, 16]. Cosmological coupling is also in agreement with the recent DESI results on the redshift evolution of the dark energy equation of state [17]. The 2023 tentative observation of cosmological coupling [6] caused a lively debate in both the observational and theoretical communities during the past year ([18, 19, 20, 21, 22] and references therein), as the arguments supporting it seemed unconvincing to some of its critics - here we provide more convincing ones.
@@ II Static black hole event horizon in a time-dependent metric
as can be read off the angular part of the line element (2.1). The equation locating the apparent horizons [^3] This equation generalizes the well known relation locating the Schwarzschild event horizon $g^{RR}=\nabla^{c}R\nabla_{c}R=1-2m/R=0$ to any spherical geometry. In spherical symmetry, the condition $\nabla^{c}R\nabla_{c}R=0$ is equivalent to the standard definition of apparent horizon in terms of the expansion scalars of ingoing and outgoing null geodesic congruences [35, 28]. (if they exist) is $\nabla^{c}R\nabla_{c}R=0$. A single root corresponds to a black hole or white hole horizon, a double root to a wormhole [28]. This equation locates also event horizons, and gives
@@
At a fixed instant of time $t=$ const., the event horizon is a 2-dimensional surface, a 2-sphere $R=$ const. spanned by the angular coordinates $\left(\vartheta,\varphi\right)$. The two-dimensional metric induced on the surface $R=$ const. is static but the time-dependent geometry just outside the horizon has $T=T(t,r)$, $a=a(t,r)$ and the areal radius $R(t,r)=a(t,r)r$ varies in time and cannot match $R=$ const. on the would-be horizon. Therefore, the junction conditions [36, 37] cannot be satisfied there: the discontinuity in the metric $g_{ab}$ induced on this 2-surface cannot be handled with distributional techniques and the jump in the metric $g_{ab}$, unlike a jump in its transversal derivative (interpreted as due to a matter layer on a null shell) has no physical interpretation [37].
@@ III Radial geodesics and Ricci scalar at the horizon
Let us examine outgoing and ingoing radial null and timelike geodesics. We will show that they cannot be continued inside the would-be static black hole horizon, which is unphysical. Furthermore, the Ricci scalar of the geometry (2.1) diverges there. The difficulties arise from assuming an exactly static horizon in a time-dependent geometry, showing that this assumption is unphysical.
@@ III.1 Radial null geodesics
III.2 Radial timelike geodesics
We now turn to timelike geodesics. The 4-velocity $u^{a}=dx^{a}/d\tau$ (where $\tau$ is the proper time) tangent to a radial timelike geodesic has components $\left(u^{0},u^{1},0,0\right)$ and is normalized to $u^{c}u_{c}=-1$, which gives
$T^{2}\left(u^{0}\right)^{2}=a^{2}\left(u^{1}\right)^{2}+1\,.$ (3.11)
Since $(u^{0})^{2}=\frac{1+(au^{1})^{2}}{T^{2}}\geq\frac{1}{T^{2}}\to+\infty$ as $x\to 0$, $u^{0}$ diverges as least as fast as $T^{-1}=2/x$ at the static horizon. The time and radial components of the geodesic equation are (using $\dot{a},T\to 0$)
$\frac{du^{0}}{d\tau}\pm\frac{2u^{0}}{a}\,\sqrt{T^{2}(u^{0})^{2}-1}\,\frac{T^{\prime}}{T}=0\,,$ (3.12)
$\frac{du^{1}}{d\tau}\pm\frac{1+a^{2}(u^{1})^{2}}{a^{2}}\,\frac{T^{\prime}}{T}=\frac{a^{\prime}}{a^{3}}\,,$ (3.13)
where, in (3.13), we used the normalization (3.11) $u^{0}=\sqrt{1+\left(au^{1}\right)^{2}}/T$. Near the horizon, Eq. (3.13) becomes
$\frac{du^{1}/d\tau}{(u^{1})^{2}+1/a^{2}}=-\frac{1}{T}\,,$ (3.14)
which can be integrated at the horizon where $T$ is static, obtaining
$u^{1}\simeq\frac{1}{a}\tan\left(\frac{\tau_{0}-\tau}{aT}\right)\,,$ (3.15)
where $\tau_{0}$ is an integration constant. As $x\to 0$, $u^{1}$ is only defined for a finite range of the proper time $\tau$. Radially infalling objects stop at the horizon and cannot enter it.
III.3 Ricci scalar
@@ VI Conclusions
The prospects of embedding an exactly stationary black hole event horizon in a FLRW (or any time-dependent) “background” are even more dire than for a spherical static horizon. In fact, a FLRW universe is not stationary and matching a horizon with axisymmetry (or any symmetry) with another metric that does not share this symmetry is doomed to failure. This fact is well known when trying to generalize the Swiss-cheese model (a Schwarzschild region matched to spherically symmetric FLRW on a 2-sphere) to axisymmetry. Matching a Kerr region to FLRW fails because the interior and exterior regions do not share axisymmetry [43, 44, 45]. This consideration, which will be discussed further in the future, extends the physical relevance of the result presented here.
Finally, we have not determined the precise way in which black hole event horizons embedded in given FLRW universes evolve with time. This law would be most interesting for astrophysics to decide whether existing puzzles can be solved by cosmological coupling, and will be addressed in the future.
@@ Acknowledgements.
We are grateful to Kevin Croker and Duncan Farrah for discussions. V.F. is supported by the Natural Sciences & Engineering Research Council of Canada (grant 2023-03234).